OpenStudy (anonymous):

Jade hiked 1 1/4 mile on the trail. Gretta hiked 1 1/2 times as far on the trail as Jade. How far did Gretta hike on the trail?

OpenStudy (wolf1728):

Greta hiked 1.5 times further than Jade so Greta's distance = 1 1/4 * 1.5 = 1.875 miles or 1 and 7/8 miles

OpenStudy (anonymous):

can you just give meh simple answer

OpenStudy (wolf1728):

Well, if you do it strictly by fractions then Greta's distance = 5/4 * 3/2 = 15/8 =1 and 7/8

OpenStudy (anonymous):

oik